### v1

# WonderHuman: Hallucinating Unseen Parts in Dynamic 3D Human Reconstruction

Abstract
In this paper, we present WonderHuman to reconstruct dynamic human avatars from a monocular video for high-fidelity novel view synthesis. Previous dynamic human avatar reconstruction methods typically require the input video to have full coverage of the observed human body. However, in daily practice, one typically has access to limited viewpoints, such as monocular front-view videos, making it a cumbersome task for previous methods to reconstruct the unseen parts of the human avatar. To tackle the issue, we present WonderHuman, which leverages 2D generative diffusion model priors to achieve high-quality, photorealistic reconstructions of dynamic human avatars from monocular videos, including accurate rendering of unseen body parts. Our approach introduces a Dual-Space Optimization technique, applying Score Distillation Sampling (SDS) in both canonical and observation spaces to ensure visual consistency and enhance realism in dynamic human reconstruction. Additionally, we present a View Selection strategy and Pose Feature Injection to enforce the consistency between SDS predictions and observed data, ensuring pose-dependent effects and higher fidelity in the reconstructed avatar. In the experiments, our method achieves SOTA performance in producing photorealistic renderings from the given monocular video, particularly for those challenging unseen parts. The project page and source code can be found at https://wyiguanw.github.io/WonderHuman/.
Index Terms:
Monocular Video, 3D Gaussian Splatting, Human Unseen Part Reconstruction, Diffusion, Score Distillation Sampling.
1 Introduction
Virtual avatars have been a key focus in computer vision, graphics, and VR/AR technologies due to their wide applications such as gaming, entertainment, communication, and telepresence. However, reconstructing high-fidelity avatars that faithfully represent human appearance, shape, and dynamics remains a formidable challenge, particularly when confronted with ubiquitous monocular video with highly limited viewpoints.
Existing avatar reconstruction methods have difficulty in reconstructing unseen parts of the human body. Previous methods [1, 2, 3, 4] typically rely on dense, synchronized multi-view inputs for the avatar reconstruction task. Recent advancements in implicit neural radians fields [5, 1, 3, 4] and 3D Gaussian Splatting [6, 7, 8, 9] have explored the high-fidelity reconstruction of both geometry and appearance of dynamic human bodies from relatively sparse multi-view videos. To reconstruct from monocular videos, other recent methods [10, 11, 12, 13, 14, 15, 16] reconstruct dynamic avatars by animating them within a canonical space derived from observation spaces using video frames. These works enable learning the inter-frame deformation to reconstruct a completed human avatar from the monocular videos. However, these methods still require the video to have full-view coverage of the human body, and typically fail to reconstruct unseen parts in the monocular video. Unfortunately, one often only has access to partial-view videos with limited viewpoints, such as front-view videos, leaving most parts of the human body unseen. Reconstructing these occluded parts thus poses a significant challenge for current methodologies.
To address this challenge, we introduce WonderHuman to achieve high-quality avatar reconstruction from partial-view monocular videos. The key idea of WonderHuman is to hallucinate the unseen parts of the human using the generative prior encoded by large-scale image diffusion models such as Zero123 [17]. The hallucinations are then combined with a Gaussian Splatting [6]-based dynamic human reconstruction framework to get a full-body avatar.
However, combining diffusion-generative priors in dynamic human reconstruction is not a trivial task with two outstanding challenges. First, the existing image diffusion generative models are designed mainly to produce single-view static images. Thus, maintaining visually accurate generated content and consistency across frames for dynamic human bodies using these generative priors is challenging; For instance, unrealistic artifacts such as blurs often appear when animating the generated bodies. Some existing works [18, 19, 20] can produce human bodies from single-view images using diffusion models, but they fail to handle dynamic cases (See Appendix B.1 for more details). Second, it’s challenging to ensure that the occluded or invisible portions of the human body generated by diffusion models are consistent with the observed visible parts. Any inconsistency between these generated and visible segments can significantly deteriorate the rendering quality of the human avatar, leading to visually incoherent results.
To tackle these issues, we present WonderHuman for high-quality dynamic human reconstruction from monocular videos. We propose leveraging the generative priors embedded in a 2D diffusion model, trained on condensed images, to infer the unseen parts of the 3D human through distillation during reconstruction. We further introduce a novel Dual-Space Optimization method to ensure visual plausibility and consistency for dynamic human representations. Our Dual-Space Optimization utilizes Score Distillation Sampling (SDS) [21] in both the canonical and observation spaces. This approach ensures that the generated content remains natural and complete by accounting not only for the information in a canonical pose but also for the dynamics across poses in the observation space. This significantly enhances the rendering quality when animating the reconstructed human avatar. Moreover, a view selection strategy and a pose feature injection approach are employed to reconcile conflicts between the SDS predictions and the given information and fuse pose-dependent effects, enhancing dynamic synthesis and overall avatar fidelity.
We conduct extensive experiments to validate the effectiveness of our method across broad benchmarks including ZJU-Mocap dataset [1], Monocap dataset [2], MVHumanNet [22] and In-the-wild dataset [23]. Compared to state-of-the-art methods [10, 15, 24, 25, 26], WonderHuman produces higher-quality photorealistic renderings of reconstructed human avatars, particularly in rendering visually plausible content for previously unseen parts of the human body. To summarize, our contributions are as follows:
  • We propose a novel framework named WonderHuman that leverages 2D generative diffusion priors to achieve high-quality, photorealistic reconstruction of dynamic humans from monocular videos, including accurate rendering of unseen body parts.
  • We introduce Dual-Space Optimization to ensure visual consistency and enhance realism throughout the dynamic reconstruction process.
  • We present a view selection strategy alongside pose feature injection to resolve conflicts between SDS predictions and observed data, ensuring pose-dependent effects and higher fidelity in the reconstructed avatar.
2 Related work
2.1 Video-based Human Avatar Reconstruction
Recently, video-based avatar reconstruction methods primarily rely on regression-based approaches [27, 28, 29, 30, 31, 32, 33, 34] or the explicit tracking of human bodies [35, 36, 37, 38, 39, 40, 41]. Since the prosperity of Neural Rendering [5], many works [42, 43, 44, 1, 2, 4, 45, 46, 47, 11, 3, 12, 10, 48, 13] try to combine neural representations with human reconstructions. These methods associate implicit neural fields on human templates like SMPL [49]. While neural representations have strong representation ability, they are slow in training. But other works [50, 14, 51, 15] additionally introduced explicit representations like meshes [50, 14], and points [51] to improve its efficiency. Yet, achieving high-quality reconstruction results using neural radiance fields still requires neural networks that are expensive to train and render. Recently, Gaussian Splatting [6] has emerged as a prominent technology, as it efficiently represents and renders complex scenes with reduced training time, without compromising quality for speed. Many recent works [7, 52, 8, 9, 53, 54, 55, 16, 56, 57, 58, 59, 60, 61, 62] try to combine Gaussian Splatting in the avatar reconstruction, which allows efficient avatar rendering in real-time. In this paper, we focus on reconstructing human avatars from monocular video. In GaussianAvatar [16], 3D Gaussians are integrated with SMPL [49] to explicitly represent humans in various poses and clothing styles. SplattingAvatar [24] embeds Gaussians onto human triangle meshes, forming a hybrid representation that significantly enhances rendering speed. Furthermore, ExAvatar [25] extends this representation to reconstruct animatable hand poses and facial expressions. However, those methods require the input video to have full-view coverage of the human body and failed to generate unseen parts in the monocular video.
2.2 Diffusion Models for Human Avatars
Pioneer works in avatar generation [63] resort to generate avatars from CLIP features [64]. Recently, diffusion models [65] show strong ability in learning complex data distributions for data generation. Some works [66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80] directly extend the SDS loss [21] to generate human avatars from text prompts. MVHuman [81] extends this framework to generate human avatars through multiview diffusion, while HumanNorm [82] integrates it with normal map generation. Additionally, HumanNorm [83] directly enables 3D human generation, benefiting from tri-plane features. Some other works [19, 84, 85, 86, 20, 18, 87, 88, 89] generate a completed human avatar from a single-view image using diffusion models. While these single-view avatar generation techniques produce avatars from single images, directly extending them to generate dynamic humans from monocular videos results in poor rendering quality for dynamic human actions. In contrast, our approach leverages the SDS loss to inpaint the unseen parts of the dynamic human body from a monocular video, with careful consideration of time coherence, consistency, and dynamics.
3 Preliminaries
3.1 3D Gaussian Splatting
3D Gaussian splatting [6] is an explicit scene representation that allows high-quality real-time rendering. The given scene is represented by a set of static 3D Gaussians, which are parameterized as follows: Gaussian center $x\in{\mathbb{R}^{3}}$, color $c\in{\mathbb{R}^{3}}$, opacity $\alpha\in{\mathbb{R}}$, spatial rotation in the form of quaternion $q\in{\mathbb{R}^{4}}$, and scaling factor $s\in{\mathbb{R}^{3}}$. Given these properties, the rendering process is represented as:
$I=Splatting(x,c,s,\alpha,q,r),$ (1)
where $I$ is the rendered image, $r$ is a set of query rays crossing the scene, and $Splatting(\cdot)$ is a differentiable rendering process. We refer readers to Kerbl et al.’s paper [6] for the details of Gaussian splatting.
3.2 Score Distillation Sampling
Score Distillation Sampling (SDS) [21] builds a bridge between diffusion models and 3D representations. In SDS, the noised input is denoised in one time-step, and the difference between added noise and predicted noise is considered SDS loss, expressed as:
$\mathcal{L}_{\text{SDS}}(I_{\Phi})\triangleq\mathbb{E}_{t,\epsilon}\left[w(t)\left(\epsilon_{\phi}(z_{t},y,t)-\epsilon\right)\frac{\partial I_{\Phi}}{\partial\Phi}\right],$ (2)
where the input $I_{\Phi}$ represents a rendered image from a 3D representation, such as 3D Gaussians, with optimizable parameters $\Phi$. $\epsilon_{\phi}$ corresponds to the predicted noise of diffusion networks, which is produced by incorporating the noise image $z_{t}$ as input and conditioning it with a text or image $y$ at timestep $t$. The noise image $z_{t}$ is derived by introducing noise $\epsilon$ into $I_{\Phi}$ at timestep $t$. The loss is weighted by the diffusion scheduler $w(t)$.
4 Method
Figure 1: Overview of WonderHuman. (1) In stage I, we reconstruct 3D Gaussians and appearances for visible human parts from partial-view videos. We start with optimizable feature vectors named canonical features capturing human geometry and appearance in a canonical space. Then, we use a Gaussian Decoder to predict Gaussian parameters and combine the Linear Blend Skinning (LBS) function with the Gaussian Splatting to render the dynamic 3D human in the observation space. (2) In Stage II, we hallucinate the invisible parts of the avatar using a Dual-space Optimization technique. We render images of the human avatar from various novel viewpoints and apply an SDS loss to learn the unseen appearances. Additionally, a normal predictor is utilized to generate normal maps that guide geometry reconstruction, while View Selection and Pose Feature Injection strategies are employed to ensure consistent appearance fusion.
Given a monocular video as the input, our goal is to reconstruct a high-quality animatable 3D human avatar including both visible and invisible parts. In WonderHuman, we employ a dynamic 3D human Gaussian representation, equipped with a generative diffusion model as hallucination prior, which produces a controllable 3D human avatar viewable from any angle. An overview of our method can be found in Fig. 1.
4.1 Stage I: Visible Appearance Reconstruction
4.1.1 Prediction of Gaussian Parameters
In the first stage, we reconstruct the visible geometry and appearance of an animatable human avatar from a partial-view monocular video. To achieve detailed and high-fidelity reconstructions, building on GaussianAvatar [16], we propose integrating normal information into the Gaussian decoder [16]. This improved decoder is used to establish a functional mapping from the underlying geometry of the human to various attributes of 3D Gaussians. And those Gaussians are initialized on the surfaces of SMPL [49] body in canonical space. Then, we have:
$(\Delta x,\Delta n,c,s)=G_{\theta}([S,S]),$ (3)
where $\theta$ represents optimizable parameters for the Gaussian decoder $G_{\theta}$, and $S$ represents the features in the canonical space. The canonical feature $S$ is an optimizable tensor, randomly initialized and optimized during training to capture texture and geometry features in canonical space. The size of $S$ is $(128\times 128)$, and it is concatenated with itself as input of Gaussian decoder $G$. This ensures that the input channel of $G$ remains $(2\times 128\times 128)$ during pose feature injection in Stage II (Sec. 4.2.3). This decoder $G$ predicts 3D center offset $\Delta{x}$, along with color and scale factors, denoted as $c$ and $s$ respectively. Additionally, it predicts normal offset $\Delta n$ that is applied to the initial SMPL normals, to capture the intrinsic geometric details. We set the opacity $\alpha$ and 3D rotation $q$ are set to fixed values of $1$ and $(1,0,0,0)$ respectively, to make the network focus more on the geometry information.
4.1.2 Dynamic Human Rendering
To render the avatar in observation space, we seamlessly combine the Linear Blend Skinning function with the Gaussian Splatting [6] process to deform the avatar from canonical space to observation space:
$I_{rgb}=Splatting(x_{o},c,Q,r),$ (4)
$x_{o}=T_{lbs}(x_{c},p,w),$ (5)
where $I_{rgb}$ represents the final rendered image. The final canonical Gaussian position $x_{c}$ is the sum of the initial position $x$ and the predicted offset $\Delta x$. The LBS function $T_{lbs}$ applies the SMPL skeleton pose $p$ and blending weights $w$ to deform $x_{c}$ into observation space as $x_{o}$, where $w$ is provided by SMPL [49]. $Q$ here denotes the remaining parameters of the Gaussians, including scale $s$, opacity $\alpha$, and rotation $q$. For more details on canonical initialization, see Appendix A.1.1.
4.1.3 Normal Map Rendering of Seen View
We aim to faithfully capture the detailed surface geometry of dynamic human bodies from partial-view videos. Central to this process is the rendering of predicted normal maps, where the predicted $\Delta n$ is applied to the initial SMPL normals $n$ to compute $n_{c}$ in canonical space. $n_{c}$ is then transformed into the observation space $n_{o}$ and rendered as normal maps $I_{n}$. The Eq. (5)&(4) are modified for normals as:
$I_{n}=Splatting(x_{o},n_{o},Q,r),$ (6)
$n_{o}=T_{lbs}(n_{c},p,w).$ (7)
This transformation maps 3D Gaussians from the canonical space to the observation space, enabling the preservation of detailed geometry encoded by the normals and appearance.
We supervise the normal vectors using high-quality normal maps derived from ground truth RGB images. For this purpose, we leverage Sapiens [90] as a normal predictor to predict normal maps from video frames, using them as supervision for normal maps rendered in our observation space, expressed as:
$\mathcal{L}_{n}=MSE(I_{n},I^{\text{gt}}_{n}),$ (8)
where $I^{\text{gt}}_{n}$ denotes the predicted normal map from Sapiens. The normal loss $\mathcal{L}_{n}$ is defined as the MSE loss between $I_{n}$ and $I^{\text{gt}}_{n}$. By aligning the predicted normal maps with those renderings, we achieve a high-fidelity representation of surface geometry that accurately captures both global and fine-grained details.
4.2 Stage II: Invisible Appearance Reconstruction
Stage I produces an animatable 3D human model with visible appearances learned from partial-view video data, but the unobserved regions of the body typically suffer from relatively low visual quality. To ensure multi-view consistency for the unseen parts, we introduce a viewpoint-conditioned diffusion model as supervision, leveraging generative priors to predict the unseen views from the given inputs. Subsequently, we optimize the Gaussian decoder $G_{\theta}$ to reconstruct a fully renderable 3D human model from any viewpoint. To effectively utilize the observations and improve consistency between observed and hallucinated results, we introduce Dual-space Optimization, View Selection and Pose Feature Injection techniques in the following.
Figure 2: Left side: Dual-pace Optimization (a) w/o dual space optimization; (b) w/ canonical optimization only; (c) w/ dual-space optimization; Right side: Pose Feature Injection (d) ground truth; (e) w/o pose feature injection; (f) w/ pose feature injection.
4.2.1 Dual-space Optimization
Zero123 [17], a viewpoint-conditioned diffusion model, is used to hallucinate full-body views from partial video frames, using reference frames from the monocular video and target view camera parameters as conditioning inputs. Its explicit view control enables precise multi-view predictions for 3D reconstruction. We leverage Score Distillation Sampling (SDS) [21] loss for predicting the unseen parts of our 3D Gaussian human model in the observation space. Unfortunately, naively combining Zero123 using SDS for dynamic human reconstruction leads to unrealistic reconstruction results. For instance, directly applying SDS in canonical space often results in degenerated quality issues in avatars—when generating 3D models with 2D diffusion models (See Fig. 2(b)).
To address this, we introduce Dual-Space Optimization, which performs SDS optimization in both canonical and observation spaces. When conducting optimization in the canonical space, we use the rendering in the canonical space from Stage I as a conditioning reference for the 2D generative diffusion model. When conducting optimization in the observation space, we utilize the selected input images from the partial-view video as conditioning references.
The SDS optimization process, combining Zero123 with the dual-space strategy, is thus expressed as:
$\mathcal{L}_{SDS}(I_{\theta})\triangleq\mathbb{E}_{t,\epsilon}[w(t)(\epsilon_{\phi}(z_{t},y,R,T,t)-\epsilon)\frac{\partial I_{\theta}}{\partial\theta}],$ (9)
where $I_{\theta}$ represents a generated image from an unseen view in observation or canonical space. $\epsilon_{\phi}$ is the predicted noise by Zero123 conditioned on the image $y$ and the target view camera parameters $(R,T)$.
Since $\mathcal{L}_{SDS}$ is applied in both canonical and observation spaces, we take the observed frames from the input video as $y_{image}$ when optimizing in observation space, and take the canonical rendering from Stage I as $y_{image}$ when optimizing in canonical space. This approach allows us to more effectively associate features across frames for the reconstruction of unseen parts.
During dual-space optimization, we found that appropriately balancing the training processes of the two spaces improves performance. As mentioned earlier, diffusion models face degeneration issues when optimizing in the canonical space. For instance, they struggle to predict accurate appearances for complex human poses in the observation space, often producing unrealistic ’tattoo-like’ appearances, as shown in Fig. 6(c). To address this, we set the weight between canonical and observation optimization as a hyperparameter in Stage II to enhances the overall process performance. This refined balance ensures better alignment of the model with the desired objectives, leading to more accurate and reliable outcomes in Stage II.
4.2.2 View Selection
The aforementioned Dual-Space Optimization with SDS aids in synthesizing the unseen appearance of human avatars. Next, we introduce view selection to analyze which regions of the avatar are poorly observed.
In both canonical and observation space optimizations, we identify the invisible views that require refinement. By utilizing the differentiable rasterization of Gaussian Splatting [6], we determine the first intersecting Gaussian for each ray, marking these as visible points. Subsequently, visibility maps are rendered to differentiate between the visible and invisible regions of the human avatar as defined in Stage I. Specificly, we first estimate the visibility of each Gaussain. During the training of Stage I with seen views, given a ray $r$, the first Gaussian hit by the ray, $x$, is marked as a seen Gaussian, and its visibility $\psi$ is set to 1. Formally:
$\psi(x,r)=\begin{cases}1,&x\text{ is the first Gaussian on }r\\ 0,&\text{otherwise}\end{cases},$ (10)
where $\psi(x,r)=1$ indicates that the Gaussian $x$ is visible in Stage I, while $\psi(x,r)=0$ represents the opposite. After Stage I training is completed, a visibility map $I_{v}$ of a random viewpoint $v$ is rendered given $\psi(x,r)$, $r$, and the remaining attributes $Q$. In $I_{v}$, if the visible region $VR(I_{v})$ covers less than 50% of the foreground region $FR(I_{v})$, this viewpoint is marked for refinement in Stage II. Then the view selection is expressed as:
$I_{v}=Splatting\left(x,\psi(x,r),Q,r\right),$ (11)
$Visibility(I_{v})=\frac{VR(I_{v})}{FR(I_{v})},$ (12)
$\mathbb{I}_{v}=\begin{cases}0,&Visibility(I_{v})\leq 50\%\\ 1,&\text{otherwise}\end{cases},$ (13)
where $\mathbb{I}_{v}=0$ signifies an unseen viewpoint, indicating that the avatar needs to be refined from viewpoint $v$ in Stage II.
Figure 3: View Selection based on visibility map (a) Seen view: Visible region (orange) covers more than 50% of the foreground region; (b) Unseen view: Invisible region (blue) covers more than 50% of the foreground region.
4.2.3 Pose Feature Injection
Furthermore, during dual-space optimization, while SDS is applied across diverse poses in observation spaces, the Gaussian decoder is trained in the canonical space. To capture pose-dependent appearances in the observation space, such as garment wrinkles in Fig. 2(d), we leverage the pose encoder similar to GaussianAvatar [16] to extract pose-related features, which are then injected into the decoder network. Consequently, we have:
$(\Delta x,\Delta n,c,s)=G_{\theta}([S,P]),$ (14)
$P=Encoder(P_{uv}),$ (15)
where $P_{uv}$ is the UV positional map of SMPL for each pose, and $P$ denotes the extracted pose feature, which is concatenated with the canonical features $S$ as input of the Gaussian decoder $G_{\theta}$. And $Encoder(.)$ maps $P_{uv}$ to $P$. All the outputs of $G_{\theta}$ remain the same as in Stage I.
4.2.4 Normal Map Supervision of Unseen View
For the reconstruction of unseen-view geometry, we extend the rendering process described in Sec. 4.1.3 to generate normal maps for unseen views. Specifically, the normal maps of given views are treated as front normal maps. To compute the back normal maps from their corresponding front normal map, we utilize a depth-aware, silhouette-consistent bilateral normal integration (d-BiNI) method [91]. These back normal maps are then combined with pretrained SMPL-aware IF-Nets [92], which inpaint the geometry of the remaining body regions. The resulting output is a complete set of normal maps, which serves as full-body normal supervision in Eq. 8.
4.3 Training Losses
In Stage I, we are modeling a dynamic avatar from partial-view videos using a Gaussian decoder. Additionally, we refine the input pose to correct inaccuracies from SMPL fitting. This stage utilizes MSE loss, SSIM loss [93], and perceptual LPIPS loss [94] between the predicted RGB images and ground truth, as $\mathcal{L}_{rgb}$, $\mathcal{L}_{ssim}$, and $\mathcal{L}_{lpips}$, respectively. We also apply Frobenius Norm loss as regularization terms for optimizable canonical features $S$, offset $\Delta x$, and scale $s$:
$\mathcal{L}_{f}^{S}=\sqrt{\sum_{k=1}^{n}|S_{k}|^{2}},\mathcal{L}_{f}^{\Delta x}=\sqrt{\sum_{k=1}^{n}|\Delta x_{k}|^{2}},\mathcal{L}_{f}^{s}=\sqrt{\sum_{k=1}^{n}|s_{k}|^{2}},$ (16)
where $\mathcal{L}_{f}^{S},\mathcal{L}_{f}^{\Delta x},\mathcal{L}_{f}^{s}$ denotes the loss of the $S$, $\Delta x$, and $s$, respectively. Combining with the normal loss $\mathcal{L}_{n}$ from Eq. 8, the total loss function for Stage I is as follows:
$\mathcal{L}_{StageI}=$ $\lambda_{rgb}\mathcal{L}_{rgb}+\lambda_{n}\mathcal{L}_{n}+\lambda_{ssim}\mathcal{L}_{ssim}$ (17)
$+\lambda_{lpips}\mathcal{L}_{lpips}+\lambda_{\Delta x}\mathcal{L}_{f}^{\Delta x}+\lambda_{s}\mathcal{L}_{f}^{s}+\lambda_{S}\mathcal{L}_{f}^{S}.$
In Stage II, the pose encoder and Gaussian decoder are optimized using SDS losses. To prevent degradation of visible appearance and geometry, $\mathcal{L}_{StageI}$ is incorporated. Additionally, $\mathcal{L}_{f}^{p}$ is added with Frobenius Norm loss to regularize the pose feature map. The total loss function for Stage II is expressed as:
$\mathcal{L}_{StageII}=$ $\mathcal{L}_{StageI}+\lambda_{p}\mathcal{L}_{f}^{p}$ (18)
$+\lambda_{SDS}(\mathcal{L}_{SDS}^{o}+\mathcal{L}_{SDS}^{c}),$
where $\mathcal{L}_{SDS}^{o}$ and $\mathcal{L}_{SDS}^{c}$ represent the SDS loss for observation space and canonical space, respectively, as defined in Eq. (9).
Furthermore, we design a progressive training strategy in this stage, gradually diminishing the weight of SDS loss. This strategy is employed to enhance further the effectiveness and efficiency of the visible appearance reconstruction. More details on progressive training are in Appendix A.2.2.
Figure 4: Qualitative comparison on four datasets. We compare the novel view synthesis quality with HumanNeRF [10], Instant-NVR [15], SplattingAvatar [24], ExAvatar [25] and GaussianAvatar [16].
Figure 5: Qualitative comparison on three datasets. We compare the novel view synthesis quality with GuessTheUnseen [26].
ZJU-Mocap(revised) MVhumanNets Monocap
[…]
HumanNeRF 19.21 0.9456 0.0715 19.38 0.9416 0.0706
Instant-NVR 19.90 0.9458 0.0630
SplattingAvatar 19.37 0.9436 0.0702 19.49 0.9467 0.0689 19.51 0.9444 0.0697
EXAvatar 19.65 0.9449 0.0678 19.69 0.9470 0.0671 19.65 0.9475 0.0676
GaussianAvatar 19.50 0.9434 0.0687 19.70 0.9471 0.0645 19.67 0.9489 0.0635
GuessTheUnseen 20.06 0.9493 0.0615 20.56 0.9502 0.0598
Ours 20.82 0.9552 0.0569 20.98 0.9517 0.0553 21.16 0.9532 0.0549
TABLE I: Quantitative evaluation on ZJU-Mocap(revised), MVHumanNet, and Monocap datasets (unseen view only).
5 Experiments
5.1 Datasets
ZJU-Mocap(revised) dataset [1]. This dataset is a multi-view dataset. We train and test using this dataset following Instant-NVR [15]. One specific camera capture is used as monocular training input and six cameras, evenly distributed around the object, are reserved for a comprehensive evaluation.
Monocap dataset. Similar to ZJU-Mocap(revised), the Monocap dataset contains multi-view videos collected by AnimatableNeRF [2] from the DeepCap dataset [35] and the DynaCap dataset [95]. The dataset setting follows the ZJU-Mocap(revised) dataset.
MVHumanNet dataset [22]. The dataset is a large-scale collection of multi-view human images, encompassing human masks, camera parameters, 2D and 3D keypoints, SMPL/SMPLX parameters. The dataset setting follows the ZJU-Mocap(revised) dataset as well.
In-the-wild dataset. This dataset contains YouTube videos collected by HumanNeRF [10] and Dance Dance Generation [23]. We employ BEV [96] to estimate camera parameters and the SMPL bodies, then utilize Xmem [97] along with Segment-anything [98] to extract foreground segmentation of video frames.
5.2 Implementation Details
All videos from all datasets are clipped to 3-5 seconds (100-150 frames) and exclusively capture front views of the subjects. During stage I, training is conducted on a single RTX-3090 GPU with a batch size of 2, requiring approximately 1 hour for 200 training epochs. In Stage II, the entire framework is trained on two RTX-3090 GPUs with a batch size of 1, while the diffusion model is loaded exclusively on the second GPU. The training process requires approximately 2–3 hours for 400 epochs, depending on the resolution.
5.3 Comparisons with Video-based Methods
We conduct comparisons of our method with HumanNeRF [10], Instant-NVR [15], SplattingAvatar [24], ExAvatar [25], GaussianAvatar [16], and GuessTheUnseen [26].
For a fair comparison, Instant-NVR [15] is trained on the revised version of the ZJU-Mocap dataset, which offers refined camera parameters, SMPL fittings, and more accurate instance masks with body-part segmentation, crucial for the execution of their method. However, HumanNeRF is not adapted to this dataset, and MVHumanNet and Monocap are applied to evaluate this method. Additionally, Instant-NVR lacks a pose refinement technique akin to HumanNeRF, which assists in addressing inaccurate fitting issues in the in-the-wild dataset. GuessTheUnseen is evaluated on the ZJU-Mocap(revised) and Monocap datasets but not on MVHumanNet, as the original images in this dataset contain black bounding boxes that significantly hinder the human motion detection process performed by GuessTheUnseen. Therefore, we will discuss the comparison results based on the type of dataset utilized.
[…]
HumanNeRF 30.23 0.9756 0.0314
SplattingAvatar 28.28 0.9693 0.0286
ExAvatar 29.46 0.9709 0.0253
GaussianAvatar 29.96 0.9716 0.0220
GuessTheUnseen 29.35 0.9685 0.0256
Ours 29.76 0.9712 0.0222
TABLE II: Quantitative evaluation on In-the-wild datasets (seen view only)
5.3.1 ZJU-Mocap(revised), MVHumanNet, and Monocap Datasets
These three datasets serve as the primary testbeds for our experiments due to the availability of ground truths for invisible parts. The quantitative results are presented in Tab. I, where our method surpasses all evaluation techniques in all three metrics, indicating its efficacy in reconstructing both geometry and appearance for invisible parts. Qualitatively, as shown in Fig. 4, the limitations of all compared methods become more evident when visualizing invisible parts. Methods such as GaussianAvatar, ExAvatar, and SplattingAvatar, which are based on Gaussian Splatting, exhibit noticeable artifacts and inconsistencies, including noisy textures and blank spots. Instant-NVR and HumanNeRF, due to their NeRF-based ray-shooting geometric reconstruction technique, not only struggle with appearance consistency but also suffer from geometric issues like floating artifacts and penetrating holes, diminishing the realism of the synthesized avatars. While GuessTheUnseen can infer unseen-view appearances, it introduces noisy textures and multi-face ‘Janus’ artifacts, as shown in Fig. 5.
[…]
full model 21.06 0.9536 0.0551
full model w/o Prog. 20.98 0.9523 0.0559
full model w/o Canonical. 20.16 0.9503 0.0586
full model w/o Opt. 19.78 0.9463 0.0626
full model w/o Observ. 19.56 0.9434 0.0685
TABLE III: Quantitative results for ablation study. Opt. includes view selection and pose feature injection. Prog. is short for progressive training strategy.
5.3.2 In-the-wild Dataset
The in-the-wild dataset comprises various monocular dancing videos sourced from the internet. For quantitative evaluation, our method demonstrates performance comparable to the baselines for the seen parts of reconstructed humans, as shown in Tab. II. However, due to the lack of novel view references, our primary focus is on qualitative evaluation results in comparison to other methods for the unseen parts of humans. As shown in Fig. 4&5, we observe that HumanNeRF faces similar challenges to Instant-NVR. The generative networks struggle to effectively fuse sampling points for novel view rendering due to a lack of supervision. This shortcoming results in floating points and “foggy” artifacts in the rendered outputs. Additionally, the Gaussian-based methods continue to produce unrealistic “tattoo-like” appearances on the backs of synthesized avatars, highlighting its limitations in preserving overall appearance fidelity. For GuessTheUnseen, the ‘Janus’ artifacts become more pronounced, and it even fails to correctly infer the appearance of some subjects. In contrast, our method consistently demonstrates superior performance in addressing the challenges of invisible parts synthesis, excelling in both geometry and appearance reconstruction. The invisible parts of the synthesized avatars show not only enhanced geometric precision but also significantly improved appearance fidelity, with fewer artifacts and smoother textures.
5.4 Ablation Study
5.4.1 Dual-space Optimization
Next, we evaluate the effectiveness of Dual-Space Optimization, with ablation results presented in Fig. 6. Observation optimization is crucial for reconstructing the invisible parts of the avatar, but it often encounters challenges and may not converge effectively, resulting in rough and less satisfactory appearances. In such cases, the canonical optimization step becomes essential. Leveraging the canonical space, the optimization converges more effectively, yielding smoother and more visually pleasing results. Nevertheless, observation optimization can mitigate the quality degradation issues that arise when relying solely on SDS optimization in the canonical space, as shown in Fig. 2(a),2(b)&2(c). This iterative approach highlights the importance of both observation and canonical optimization for achieving optimal reconstruction outcomes.
5.4.2 View Selection and Pose Feature Injection
We investigate the influence of View Selection and Pose Feature Injection in the following. As shown in Fig. 7, view selection filters out visible views, preserving the alignment between visible and canonical appearances, thereby reducing potential disruptions from observation optimization. Additionally, pose feature injection plays a crucial role in further enhancing dynamic appearance, allowing for the capture of finer details, especially in facial regions and cloth textures. These improvements significantly contribute to the overall fidelity and realism of the synthesized avatars.In Tab. III, we present a quantitative evaluation of all components in the ZJU-Mocap(revised) dataset, MVHumanNet dataset, and Monocap dataset. Our full model, coupled with progressive training, achieves the best results in this evaluation.
Figure 6: Ablation study about Dual-pace Optimization (a) conditioning image for SDS. (b) w/o dual space optimization. (c) w/ observation optimization only. (d) full model novel view.
Figure 7: Ablation study about View Selection and Pose Feature Injection (a) w/o both. (b) w/ view selection and w/o pose feature injection. (c) full model novel view.
5.5 Novel Poses Animation
Our method aligns the generated Gaussian human avatars with the SMPL model, enabling us to animate the reconstructed avatar with novel poses, as shown in Fig. 8. Please refer to the accompanying video for dynamic results.
Figure 8: Avatar animation with novel poses.
5.6 Efficiency
As shown in Fig. 1, our method, with its two-stage training process, requires approximately 3 training hours, significantly outperforming HumanNeRF, which demands about 10 hours of training on the same device. Furthermore, our method can achieve almost real-time rendering speed at 18 fps. But Instant-NVR and HumanNeRF can only render with 2 fps and 7 fps respectively.
6 Discussion and Conclusion
6.1 Limitation
Since our method depends on human body fitting and foreground segmentation, artifacts may occur due to inaccuracies in these videos within the processes. Despite incorporating pose optimization to correct poses, the reconstruction of hand parts and body shape may still exhibit artifacts in certain cases, as shown in the videos. While our approach generally yields more realistic results, similar to many existing methods [15, 10, 16], it still faces challenges in accurately modeling loose attire, such as dresses, underscoring areas for potential improvement in future iterations.
6.2 Conclusion
In this paper, we introduce WonderHuman, a novel approach for high-quality dynamic human reconstruction from monocular videos. By leveraging 2D diffusion model priors, WonderHuman effectively reconstructs and infers the unseen parts of 3D human avatars. We introduce Dual-Space Optimization, which applies Score Distillation Sampling (SDS) in both canonical and observation spaces, ensuring visual consistency and realism across various poses. Furthermore, View Selection and Pose Feature Injection strategies resolve conflicts between SDS predictions and observed data, enhancing overall avatar fidelity. Extensive experiments on benchmarks demonstrate that WonderHuman outperforms state-of-the-art methods, particularly in rendering the unseen parts of the human body.
Appendix A Training Details
This section provides more details about the implementation and training of our method.
A.1 Stage I
A.1.1 Canonical Initialization
We unwrap the T-pose body onto a UV map, where each pixel stores a 3D position vector. The positional UV map, with a resolution of $(512\times 512\times 3)$, is used to initialize Gaussians in the canonical space, ensuring proper alignment with the body’s structure. Additionally, a downsampled $(128\times 128\times 3)$ version of the positional UV map serves as input to the Gaussian decoder, aiding in reconstructing and refining the 3D representation. Furthermore, we use blend
A.1.2 Training
The training objectives in this stage focus on image losses and optimizations about Gaussian parameters.We set weights for each objective as $\lambda_{rgb}=0.8$, $\lambda_{n}=0.8$, $\lambda_{ssim}=0.2$, $\lambda_{lpips}=0.2$, $\lambda_{\Delta x}=0.85$, $\lambda_{s}=0.03$, $\lambda_{S}=1$.
A.1.3 Pose Optimization
Our method leverages pose optimization from GaussianAvatar [16] for the In-the-wild dataset as a correction for fitted SMPL [49] pose parameters. We have omitted this functionality for the ZJU-Mocap dataset, as their ground truth pose is accurate. However, GausianAvatar keeps optimizing pose parameters for ZJU-Mocap dataset, which leads to inaccurate poses, especially for invisible parts. Please check the accompanying video results for more details.
A.2 Stage II
A.2.1 Dual-space Optimization
In this stage, we apply Dual-space optimization on top of visible appearance reconstruction to predict the invisible appearance. During training, each epoch is divided into three parts: 50% for given view training and 50% for Dual-space optimization. In Dual-space optimization, the weight of canonical optimization is treated as a hyperparameter, defaulting to 50%. the The fine-tuning losses are added upon $\mathcal{L}_{StageI}$. We set $\lambda_{p}=0.5$ and $\lambda_{SDS}=0.3$ initially.
A.2.2 Progressive Training
We design a progressive training strategy in this stage, gradually diminishing the weight of SDS loss. This strategy is employed to enhance further the effectiveness and efficiency of the visible appearance reconstruction. Based on this strategy, the $\lambda_{SDS}$ is reduced gradually by following:
$\lambda_{SDS}(t)=\lambda_{SDS,0}\cdot\frac{1}{2^{\lfloor\frac{t-t_{\text{0}}}{k}\rfloor}}$ (19)
where $t$ and $t_{0}$ are the current epoch and starting epoch respectively, $k$ is the interval step of changing the weight. We set $t_{0}=100$ and $k=100$.
A.3 Resolution
The video resolution for the ZJU-Mocap (revised)[17] and Monocap datasets is consistently maintained at $1024\times 1024$ pixels, while MVHumanNet[22] has a resolution of $2048\times 1500$ pixels. For videos collected from the internet, the resolution ranges from 720p to 1080p. However, in Stage II, Zero123 only accepts $256\times 256$ as input. Therefore, for SDS loss calculation, we crop the ground truth images based on their masks and resize them to $256\times 256$.
Figure 9: Qualitative comparison results with SIFU [99] and SITH [18].
Appendix B More Experiments
B.1 Comparison with Image-based Methods
In this section, we compare our method with SIFU [99], SITH [18], and ELICIT [100], all of which are single-image reconstruction techniques designed to synthesize unseen parts of human avatars.
SIFU proposes an approach to reconstruct clothed human avatars from single images. Qualitatively, as shown in Fig. 9, this method can reconstruct decent geometry but fails to synthesize the texture of unseen parts of humans. SITH, similar to SIFU, is a method for single-image reconstruction. SITH can predict the texture of unseen parts of humans, but their generated textures contain unrealistic artifacts.
Figure 10: Qualitative comparison results with ELICIT [100] on novel poses.
[…]
MVHumanNet SIFU 19.29 0.9486 0.0706
SITH 19.68 0.9462 0.0699
Ours 20.98 0.9517 0.0553
Monocap SIFU 18.96 0.9406 0.0659
SITH 19.06 0.9428 0.0.0673
Ours 21.16 0.9532 0.0549
ZJU-Mocap(revised) ELICIT 19.23 0.9456 0.0.0689
Ours 20.82 0.9552 0.0569
TABLE IV: Quantitative evaluation on MVHumanNet, ZJU-Mocap(revised), and Monocap datasets.
ELICIT is a generative model that takes one image and a motion sequence as input to generate an animatable avatar. Qualitative results are shown in Fig. 10. For a fair comparison, since our method takes an image sequence as input, we are comparing the quality by synthesizing a novel pose that is not included in our inputs. Even though ELICIT can predict the unseen parts of humans, it shows blurred edges and floating artifacts while applying motions. Because only one image is used as input for ELICIT, the texture cannot be adapted to novel poses dynamically. In contrast, our method associates texture to different body parts across frames and can predict the correct texture for unseen parts robustly.
In Tab. IV, we present the quantitative evaluation results. SIFU and SITH were tested on the Monocap dataset, while ELICIT was evaluated on the ZJU-Mocap(revised) dataset. The results demonstrate that our method consistently achieves superior performance compared to the state-of-the-art approaches, underscoring its efficacy and robustness.